how many pounds of coffee worth $9 a pound should be added to 30 pounds of coffee worth $3 a pound to get a mixture worth $4 a pound
Answer by stanbon(75887)

how many pounds of coffee worth $9 a pound should be added to 30 pounds of coffee worth $3 a pound to get a mixture worth $4 a pound
------------
Equation:
value + value = value
9x + 3*30 = 4(x+30)
-----
9x + 90 = 4x + 120
4x = 30
x = 7.5 lbs (amt. of $9 coffee to add)
=====
Cheers,
Stan H.
==========